OpenStudy (anonymous):

A player is halfway between 1st base and 2nd base. Find the player's coordinates

OpenStudy (anonymous):

1st base is 90,0 and 2nd base is 90,90

OpenStudy (anonymous):

Half = 1/2 of Sum

OpenStudy (anonymous):

will it be 0, 45

OpenStudy (anonymous):

( 90,0 ) ( 90,90 ) (x1, y1) ( x2, y2) x = (x1 + x2) /2 y = ( y1 + y2) /2 @lvasquev462 Plug them in and Check it yourself

OpenStudy (anonymous):

after we add them i got 180 and 90 so do we square them now?

OpenStudy (anonymous):

x = (x1 + x2) /2

OpenStudy (anonymous):

Divide by 2 = 1/2 = half = midpoint :)

OpenStudy (anonymous):

?????

OpenStudy (anonymous):

can u draw it ?

OpenStudy (anonymous):

am homeschooled this is hard

OpenStudy (anonymous):

where did you get 180 , 90?

OpenStudy (anonymous):

Midpoint = average of coordinate? Do you understand average ?

OpenStudy (anonymous):

45,45 ?/ is that the anwers

OpenStudy (anonymous):

where did you get 180 , 90?

OpenStudy (anonymous):

i did it wrong sorry

OpenStudy (anonymous):

what do u get

OpenStudy (anonymous):

180 , 90 is on the right path...

OpenStudy (anonymous):

x = (x1 + x2) /2 = ( 90 + 90) /2 = 90

OpenStudy (anonymous):

Do you see the way I plug them in! Can you mimic for: y = ...

OpenStudy (anonymous):

Yes what did u get and can u draw it step by step

OpenStudy (anonymous):

I'll need @Nameless help, because I already try step by step, but you don't read at all as usual :(

OpenStudy (phi):